# Ada-NAV: Adaptive Trajectory Length-Based Sample Efficient Policy Learning for Robotic Navigation

### v4

@@ IV-A Evaluation Metrics
We use the following metrics to evaluate the quality of the navigation paths generated with a fixed sample budget:
@@
  • Average path length: The average length of the path taken to reach a goal located 10 meters away from the robot’s starting position.
@@ IV-B Results and Discussions
Policy Learning: We observe in Figure 4 and 6 that, for the outdoor robot simulator, Ada-NAV achieves similar or better reward returns with significantly fewer samples under both the navigation settings compared to the fixed and random trajectory lengths. Significantly, despite providing comparable returns to Ada-NAV, the sample efficiency of the constant and large trajectory lengths is significantly lower (see the bar plots in Fig. 4 and 6). Please see the Appendix for similar results in 2D gridworld environments with actor-critic algorithm[49].
Navigation Performance: We compare our method’s navigation performance qualitatively in Fig. 5 and quantitatively in Table I. To highlight the importance of sample efficiency during training, we use policies trained for a fixed sample budget to perform navigation tasks in even and uneven simulated terrains. We observe that, except Ada-NAV, the policies cannot consistently complete the navigation tasks, resulting in significantly lower success rates. Even successful trials of these policies lead to significantly longer paths with several loops instead of heading toward the goal. In contrast, Ada-NAV results in relatively straight and consistent paths toward the goals. We further observe that uneven terrain navigation is a comparatively challenging task even for traditional methods such as Ego-graph[48]. Hence, the success rates are relatively low due to the robot flip-overs that occurred during trials. However, we observe that Ada-NAV leads to higher success rates even compared to baseline methods such as ego-graph. We further observe that Ada-NAV results in comparatively lower elevation cost indicating that the generated paths are smoother in uneven terrain conditions. Finally, we deploy the policies trained with the outdoor simulator in the real world using a Clearpath Husky robot. We observe that our Ada-NAV even and uneven terrain navigation policies can be transferred to the real world without significant performance degradation in terms of success rate, path length, and elevation cost while other policies demonstrate inconsistent navigation performance similar to the simulations.
@@ V Conclusions, Limitations, and Future Works
In this work, we propose an adaptive trajectory scheme, called Ada-NAV, which encourages exploration and induces training sample efficiency for goal-reaching navigation in the sparse reward setting with RL-based methods. We do so by empirically establishing a positive correlation between a policy’s entropy and the spectral gap of its induced Markov Chain. We show, through experiments in navigation tasks, that Ada-NAV induces more sample-efficient training than both fixed and random trajectory length schemes without the need for tuning. In experiments on real robots, Ada-NAV-trained policies generally achieve better success rates and shorter path lengths than policies trained using constant and random trajectory lengths for the same sample budget.
An important limitation of the present work is the restricted scope of our experimental investigation of the connection between policy entropy and the induced spectral gap. We considered only simple gridworld environments for the purposes of this paper, yet this important connection merits further investigation and validation on more sophisticated challenges, such as the MuJoCo suite of RL problems [50]. During real-world experiments, we observe that all the navigation policies often lead to high motor vibrations due to dynamically infeasible velocities provided as the action. Hence, further investigation is required to generate dynamically feasible and smooth actions for a better sim-to-real transfer.

### v7

@@ IV-A Evaluation Metrics
  • Average path length: The average length of path taken to reach goal 10 meters away from the starting position.
@@ IV-B Results and Discussions
Policy Learning: We observe in Figure 4 and 6 that, for the outdoor robot simulator, CCE achieves similar or better reward returns with significantly fewer samples under both the navigation settings compared to the fixed and random trajectory lengths. Significantly, despite providing comparable returns to CCE, the sample efficiency of the constant and large trajectory lengths is significantly lower (see the bar plots in Fig. 4 and 6). For the agents using CCE, in the even terrain simulations, all agents started with an initial trajectory length of $180$, and in the uneven terrain, $200$. All those agents end with a final length of $300$.
Navigation Performance: We compare our method’s navigation performance qualitatively in Fig. 5 and quantitatively in Table I. To highlight the importance of sample efficiency during training, we use policies trained up to episode $150$ for a limited sample budget of $\sim 4.5e4$ to perform navigation tasks in even and uneven simulated terrains. The policies, except for ones from CCE, cannot consistently complete the navigation tasks, resulting in significantly lower success rates. Even successful trials of these policies lead to significantly longer paths with several loops instead of heading toward the goal. In contrast, CCE results in relatively straight and consistent paths toward the goals. We further observe that uneven terrain navigation is a comparatively challenging task even for traditional methods such as Ego-graph[51]. Hence, the success rates are relatively low due to the robot flip-overs that occurred during trials. However, CCE leads to higher success rates even compared to baseline methods such as ego-graph. We further observe that CCE results in lower elevation cost indicating that the generated paths are smoother in uneven terrain conditions. Finally, we deploy the policies trained with the outdoor simulator in the real world using a Clearpath Husky robot. With CCE, even and uneven terrain navigation policies can be transferred to the real world without significant performance degradation in terms of success rate, path length, and elevation cost while other policies show inconsistent performance similar to the simulations.
Remark: Given more training samples, the baseline methods could perform better in success rate and path length during test time given that Figures 4 and 6 show the many times they can reach similar cumulative return as CCE. However, we tested policies from earlier in training to highlight that CCE does not need as many samples.
@@ V Conclusions, Limitations, and Future Works
We propose CCE, which controls exploration and induces training sample efficiency for goal-reaching navigation in the sparse reward setting with RL-based methods by adaptively changing the trajectory length during training. We do so by empirically establishing a positive correlation between a policy’s entropy and the spectral gap of its induced Markov Chain. We show that CCE induces more sample-efficient training than baselines. On real robots, CCE-trained policies generally achieve better success rates and shorter path lengths than ones trained with constant trajectory lengths and entropy regularization for the same sample budget.
An important limitation of the present work is the restricted scope of our experimental investigation of the connection between policy entropy and the induced spectral gap. Although we cannot explore the connection in complex environments due to spectral gap calculation impracticality, it can still merit investigation for simple environments with continuous state and action spaces. During real-world experiments, all the navigation policies often lead to high motor vibrations from infeasible velocities provided as the action. Hence, further investigation is required to generate dynamically feasible and smooth actions for better sim-to-real transfer.